A platinum sphere with radius 0.0103 m is totally immersed in mercury. Find the weight of the sphere, the buoyant force acting on the sphere, and the sphere\'s apparent weight. The densities of platinum and mercury are 2.14 × 104 kg/m3 and 1.36 × 104 kg/m3, respectively.
Answer #1

r = 0.0103 m

V = Volume of sphere = 4\pir3/3 = 4 (3.14) (0.0103)3 /3 = 0.000046 m3

\rhop = density of platinum = 2.14 x 104

m = mass of platinum = \rhop V = (2.14 x 104) (0.000046) = 0.98 kg

weight = mg = 0.66 x 9.8 = 9.6 N

Buoyant force = Fb = \rhom V g = (1.36 x 104) (0.000046) (9.8) = 6.1 N

apparent weight = weight - Buoyant force = 9.6-6.1 = 3.5 N